# Perturbing the Gradient for alleviating Meta Overfitting

### v1

@@ 4 Proposed Methodology
The proposed optimisation method involves adding a noise term to the gradient direction at each step of optimisation. Due to its statistical propoerties and general usefullness, the noise is considered to be drawn from a Gaussian with a zero mean and non-trivial standard deviation. This addition of the noise term to the gradient can also be considered as incrementing the diversity of the tasks drawn from the distribution, as the gradient direction of each task is changed from its default. This diverts the model parameters from moving in a deterministic direction towards the minima given by the gradient descent process. As a result, due to perturbations of the gradient at any time step, the gradient at the next step gets altered from its otherwise supposed direction and eventually all subsequent steps are influenced. Thus, in our methodology, the effective gradient becomes probabilistic in nature and is governed by the resultant noise accumulated over all of the steps. However, to enforce convergence of the model the noise is scaled down by a factor after a certain number of iterations. This optimisation method can be employed in both the inner loop or the outer loop of MAML as well as simultaneously.
@@
To explain the working of this approach, we can assume a 3d landscape with hills and valleys that represent the meta-loss landscape over a 2d parameter space, such that a valley at its lowest point represent the minimum loss in some proximity and the corresponding co-ordinate of its location represents a locally optimal parameter. Starting from any random point in the landscape and traversing along the direction of gradient, tends to move towards the a nearby valley i.e. a point where the loss is minimum. Ideally, this parameter is considered optimal for the task at hand. However, under the condition of non-mutually exclusive task-setting, location of the deepest valley or point of minimum loss is by design, the parameters that encourages memorization. Thus, under such circumstances, the model suffers from overfitting. On the other hand, when we add a Gaussian noise to the direction of gradient descent, it diverts the trajectory from its pre-determined path and causes it to probabilistically arrive at some other nearby valley, the parameters for which may not be the most optimal in terms of training loss but which will be sufficiently optimal to perform the training tasks and at the same time prevent the model from overfitting. This in turn enables generalization to new tasks because in this case relying on a single global function does not minimize loss for the model and is hence encouraged to utilize the support set.
@@ 5 Feedback from Meta-Test to Meta-Train
Meta overfitting can also be viewed and indicated from the inability of a meta learning model to perform well on an unseen task during meta-testing but comparatively much better performance on meta-training tasks. Thus, to close the disparity between the two scenarios, we propose to change the evaluation setting and allow the newly seen task to provide a feedback to the meta-training phase. This feedback mechanism acts as crucial information for the model to adapt the meta-parameters during re-meta-training so as to escape from the valley of the overfitting parameter instance.
In the proposed feedback mechanism, the parameters initially obtained after meta-training process are further retrained to move in the direction of the optimal test task parameters. For this purpose, the model is tested on a meta-test task to evaluate its performance which tends to be suboptimal in the meta-overfitting case and simultaneously the gradient of the test task loss at the meta learnt parameters is recorded. This indicates the direction in which the meta-parameters must have essentially been for performing well on the test task. We hence leverage on this gradient direction to move the meta-parameters towards the test-task optimal parameters by virtue of retraining the meta-trained model with more meta-training tasks. For every task in the meta-training set, the outer loop gradient is weighed according to its cosine similarity with the test task gradient. Hence, the meta-training tasks that exhibit gradients similar to the test task will be given higher weightage and thereby pulling the updated meta-parameters towards the gradient direction given by the test task. This proposed method enables the model to climb out of the overfitting parameters and move towards a point in the parameter space which is more suited for the test task. Moreover, this approach is also space efficient as there is no need for storing the test task throughout the retraining and merely the gradients need to be stored for comparison. The algorithm is given in 2
@@ 6 Experiments and Results
The experiments are performed on an Intel(R) Xeon(R) CPU E5-2630 v3 processor running at 2.40GHz with 16GB RAM and Quadro K2200 Graphics controller. Throughout the experiments, we aim to address that a model can be prevented or recovered from meta-overfitting with the methods described in this paper. We examine the results of our proposed approach and make comparisons with existing methods that deal with meta-overfitting, on both regression and classification tasks. For regression problems, we consider sinusoidal regression and pose prediction tasks whereas for classification we consider the Omniglot, MiniImageNet and D’Claw datasets under a non-mutually exclusive task setting. For formulation of non-mutually exclusive settings, this paper draws inspiration from the methodology described in the experiments of Rajendran et al. (2020). The code is available at Github.
@@ 6.2 Classification on Omniglot, MiniImageNet and D’Claw
For classification, the task formulation is done in two different ways [Meta-Augmentation] based on the level of randomness or the conditional entropy : Intershuffle and Non-mutually exclusive. The intershuffle mode is the default way of task formation in few shot learning, wherein for k way classification, the sampled k classes in any task are randomly assigned a discrete label from [0, k-1] and a class may have different labels when occurring in different tasks. In case of non-mutually exclusive, the classes are divided into k-sized disjoint partitions, and while creating a task from a particular partition, labels are assigned to the task in an ordered fashion. This ascertains that any specific class is always assigned the same label.
For our experiments on the few shot classification problem, we consider three benchmark datasets Omniglot, MiniImageNet and D’Claw. Further we evaluate MAML, TAML, MR-MAML along with our approach on the above datasets under the same task settings and provide comparisons.
6.2.1 Omniglot
The Omniglot dataset [Lake et al. (2019)] comprises of 50,000 handwritten characters as 105x105 grayscale images from 50 different alphabets, including modern and ancient scripts. Multiple individuals wrote each character, producing variations in writing style, orientation, and stroke thickness. The dataset is partitioned into a "background" set of 30 alphabets with 964 characters each and an "evaluation" set of 20 alphabets with 659 characters each. No character appears in both sets, ensuring there is no overlap between the two.
For the Omniglot dataset, the tables 2 and 3 shows the results of classification in the non mutually exclusive case. For both the 20 way and 5 way setting, as compared to vanilla MAML and TAML, our proposed approach is able to improve the classification accuracy significantly and gives comparable results with MR-MAML. For both the cases, the standard deviation of the noise is set to a scale of $10^{-7}$.
@@ 7 Conclusion
This work is based on the assumption that in order to alleviate from meta-overfitting, the model should be prevented from following its conventional gradient direction in the parameter space that leads it towards a point of minimum loss. In light of the above, we proposed an approach that Experiments have shown that the proposed model is able to handle non-mutually exclusive task settings and performs better than the existing state-of-the-art methods for meta-overfitting. The benefit of our approach is that the method does not make any changes to the learning capacity of the model by regularization nor to the data, but only to the optimization procedure to reach at a different optimum altogether. Additionally, we have also introduced a mechanism for providing feedback from the meta-test to the meta-train phase in order to ensure improved performance on the test task in a meta-overfitting scenario.

### v2

@@ 5 Proposed Methodology
The proposed optimisation method involves adding a noise term to the gradient direction at each step of optimisation. Due to its statistical properties and general usefullness, the noise is considered to be drawn from a Gaussian with a zero mean and non-trivial standard deviation. This addition of the noise term to the gradient can also be considered as incrementing the diversity of the tasks drawn from the distribution, as the gradient direction of each task is changed from its default. This diverts the model parameters from moving in a deterministic direction towards the minima given by the gradient descent process. As a result, due to perturbations of the gradient at any time step, the gradient at the next step gets altered from its otherwise supposed direction and eventually all subsequent steps are influenced. Thus, in our methodology, the effective gradient becomes probabilistic in nature and is governed by the resultant noise accumulated over all of the steps. However, to enforce convergence of the model the noise is scaled down by a factor after a certain number of iterations. This optimisation method can be employed in both the inner loop or the outer loop of MAML as well as simultaneously.
@@
To explain the working of this approach, we can assume a 3d landscape with hills and valleys that represent the meta-loss landscape over a 2d parameter space, such that a valley at its lowest point represent the minimum loss in some proximity and the corresponding co-ordinate of its location represents a locally optimal parameter. Starting from any random point in the landscape and traversing along the direction of gradient, tends to move towards the a nearby valley i.e. a point where the loss is minimum. Ideally, this parameter is considered optimal for the task at hand. However, under the condition of non-mutually exclusive task-setting, location of the deepest valley or point of minimum loss is by design, the parameters that encourages memorization. Thus, under such circumstances, the model suffers from overfitting. On the other hand, when we add a Gaussian noise to the direction of gradient descent, it diverts the trajectory from its pre-determined path and causes it to probabilistically arrive at some other valley, the parameters for which may not be the most optimal in terms of training loss but which will be sufficiently optimal to perform well on the testing tasks and at the same time prevent the model from overfitting. This in turn enables generalization to new tasks because in this case relying on a single global function does not minimize loss for the model and is hence encouraged to utilize the support set.
@@ 6 Feedback from Meta-Test to Meta-Train
Meta overfitting can also be viewed and indicated from the inability of a meta learning model to perform well on an unseen task during meta-testing but comparatively much better performance on meta-training tasks. Thus, to close the disparity between the two scenarios, we propose to change the evaluation setting and allow the newly seen task to provide a feedback to the meta-training phase. This feedback mechanism acts as crucial information for the model to adapt the meta-parameters during re-meta-training so as to escape from the valley of the overfitting parameter instance. This allows for an improved performance on that particular unseen task. This is substantial because whenever the model overfits, it fails to fine-tune to any task from the meta-test set which comprises of classes that are mutually exclusive or completely unseen in the meta-train. Fine-tuning extensively further leads to overfitting to the few samples within the specific task. However, the feedback mechanism proposed here allows a model to recover and perform well on any desired task from the meta-test after fine-tuning, while maintaining generality of the samples within the task.
In the proposed feedback mechanism, the parameters initially obtained after meta-training process are further retrained to move in the direction of the optimal test task parameters. For this purpose, the model is tested on a meta-test task to evaluate its performance which tends to be suboptimal in the meta-overfitting case and simultaneously the gradient of the test task loss at the meta learnt parameters is recorded. This indicates the direction in which the meta-parameters must have essentially been for performing well on the test task. We hence leverage on this gradient direction to move the meta-parameters towards the test-task optimal parameters by virtue of retraining the meta-trained model with more meta-training tasks. For every task in the meta-training set, the outer loop gradient is weighed according to its cosine similarity with the test task gradient. Hence, the meta-training tasks that exhibit gradients similar to the test task will be given higher weightage and thereby pulling the updated meta-parameters towards the gradient direction given by the test task. This proposed method enables the model to climb out of the overfitting parameters and move towards a point in the parameter space which is more suited for the test task. Moreover, this approach is also space efficient as there is no need for storing the test task throughout the retraining and merely the gradients need to be stored for comparison. The algorithm is given in 2. Hence by using the feedback mechanism, performance on a particular task can be improved even in the case of meta-overfitting.
@@ 7 Experiments and Results
The experiments are performed on an Intel(R) Xeon(R) CPU E5-2630 v3 processor running at 2.40GHz with 16GB RAM and Quadro K2200 Graphics controller. Throughout the experiments, we aim to address that a model can be prevented or recovered from meta-overfitting with the methods described in this paper. We examine the results of our proposed approach and make comparisons with existing methods that deal with meta-overfitting, on both regression and classification tasks. For regression problems, we consider sinusoidal regression and pose prediction tasks whereas for classification we consider the Omniglot, MiniImageNet and D’Claw datasets under a non-mutually exclusive task setting. For formulation of non-mutually exclusive settings, this paper draws inspiration from the methodology described in the experiments of Rajendran et al. (2020).
7.1 Model Architecture
The base architecture of the neural network model used for both classification and regression is same as the implementation of vanilla MAML by Finn et al. (2017). The choice for MAML as the base architecture is due to its extensive study in the research community and widespread applicability. The classification model used in the experiments consists of 4 convolutional blocks followed by a fully connected classifier layer. Each convolutional block comprises of a conv2d layer with 64 filters of size 3x3, a batch normalization layer, a max pooling layer and a ReLU activation. The neural network model used for classification is composed of 2 hidden layers with 40 neurons along with ReLU used for nonlinearities. For all the experiments, the initialization parameters for meta-training are drawn from a truncated normal distribution with 0 mean and 0.01 standard deviation.
For perturbing the gradients, we augment the optimization procedure. This is implemented in a straightforward manner by adding a noise term to the usual gradient computed through automatic differentiation in the model with respect to the loss. The noise term is sampled from a normal distribution and is a hyperparmater specific to a dataset.
In the Feedback method experiments, initially the direction of the gradients for a test task is determined and stored. Now while retraining the model, all the training tasks are weighed according to the cosine similarity of its gradient and the stored direction. This gradually pulls the model parameters in the direction of the test task. The code for the experiments is available at Github.
@@ 7.3 Classification on Omniglot, MiniImageNet and D’Claw
For classification, the task formulation is done in two different ways [Meta-Augmentation] based on the level of randomness or the conditional entropy : Intershuffle and Non-mutually exclusive. The intershuffle mode is the default way of task formation in few shot learning, wherein for k way classification, the sampled k classes in any task are randomly assigned a discrete label from [0, k-1] and a class may have different labels when occurring in different tasks. In case of non-mutually exclusive, the classes are divided into k-sized disjoint partitions, and while creating a task from a particular partition, labels are assigned to the task in an ordered fashion. This ascertains that any specific class is always assigned the same label. As a result, the model gets to learn the class and label binding which it memorizes to efficiently generate lower loss during training.
For our experiments on the few shot classification problem, we consider three benchmark datasets Omniglot, MiniImageNet and D’Claw. Further we evaluate MAML, TAML, MR-MAML along with our approach on the above datasets under the same task settings and provide comparisons. The evaluation metric for all the classification tasks is considered to be accuracy, as the support and query set for all the tasks are designed to be balanced in terms of class label.
The Omniglot dataset [Lake et al. (2019)] comprises of 50,000 handwritten characters as 105x105 grayscale images from 50 different alphabets, including modern and ancient scripts. Multiple individuals wrote each character, producing variations in writing style, orientation, and stroke thickness. The dataset is partitioned into a "background" set of 30 alphabets with 964 characters each and an "evaluation" set of 20 alphabets with 659 characters each. No character appears in both sets, ensuring there is no overlap between the two. The Omniglot dataset is augmented with rotations by multiples of 90 degrees, as proposed by Santoro et al. (2016).
For the Omniglot dataset, the tables 2 and 3 shows the results of classification in the non mutually exclusive case. For both the 20 way and 5 way setting, as compared to vanilla MAML and TAML, our proposed approach is able to improve the classification accuracy significantly and gives comparable results with MR-MAML. For both the cases, the standard deviation of the noise is set to a scale of $10^{-7}$. In comparison to vanilla MAML, where the model gets stuck in a overfitting minima preventing generalisation, perturbing the gradients using our approach allows the model to lift itself from the meta-overfitting minima and converge to another minima in the parameter space.
@@ 8 Conclusion
This work is based on the assumption that in order to alleviate from meta-overfitting, the model should be prevented from following its conventional gradient direction in the parameter space that leads it towards a point of minimum loss. In light of the above, we proposed an approach that augments the optimization process to add a subtle noise from a noise distribution, in the gradient direction proposed by the vanilla optimizer. Experiments have shown that the proposed model is able to handle non-mutually exclusive task settings and performs better than the existing state-of-the-art methods for meta-overfitting. The benefit of our approach is that the method does not make any changes to the learning capacity of the model by regularization nor to the data, but only to the optimization procedure to reach at a different optimum altogether. The standard deviation of the noise distribution is observed to be a hyperparameter, a careful selection of the values of which, results in optimized performance and shows a diminishing performance towards either ends of the scale of values. Additionally, we have also introduced a mechanism for providing feedback from the meta-test to the meta-train phase in order to ensure improved performance on a test task drawn from the meta-testing classes under a meta-overfitting scenario. The feedback guides the model parameters in the direction of the test task to a minima that is closer to the specific task. This improves the performance of the model and alleviates meta-overfitting.